g(x)=(x+2)^2
____________
x^2+5x+6

Answer by ewatrrr(24785)

 
Hi
g(x) = (x+2)^2/(x^2+5x+6)
g(x) = (x+2)^2/(x+2)(x+3) = (x+2)/(x+3)